OpenStudy (anonymous):

Harold collected 160 different postage stamps. Forty-five of his stamps had pictures of presidents. medal and fan What percent of the stamps pictured presidents? Round to the nearest percent, if necessary.

OpenStudy (homework123456):

medal me and ill tell you @ambergorney

OpenStudy (anonymous):

who ever explaines the best will get medal but both of you will get fan:)

OpenStudy (anonymous):

In order to find this, we first subtract the 45 from the 160:\[160 - 45 = 115\] Then, we divide the 115 by 160:\[\frac{ 115 }{ 160 } = 0.718\] Which means 72% of them are NOT stamps with presidents on them. Since we're looking for 100%, we'd just subtract 100 by 72:\[100 - 72 = 28\] That means that 28% of the 100% of the stamps are presidential stamps.

OpenStudy (anonymous):

Make sense? @ambergorney

OpenStudy (anonymous):

ok thank you